what is the vertex form of y=3x^2-24x +21
Found 2 solutions by nerdybill, Onmyoji.S:
Answer by nerdybill(7384)

what is the vertex form of y=3x^2-24x +21
.
We can find it by "completing the square":
y=3x^2-24x +21
group x's:
y=(3x^2-24x) +21
factor out a 3:
y=3(x^2-8x) +21
complete square:
y=3(x^2-8x+16) +21 - (3*16)
y=3(x^2-8x+16) +21 - 48
y=3(x^2-8x+16) - 27
y=3(x-4)^2 - 27 (this is the vertex form)

Answer by Onmyoji.S(32)

y=3x²-24x+21
vertex form:-b/2a
= -(-24)/2*3
=4